4 1/9 buses go on a field trip, maxium compacity is 72 students. How many actually students go on the field trip?
Answer by benazir.sj@gmail.com(109)

number of buses gone for field trip =4 1/9 =37/9
number of students in 1 bus =72
total number of students gone for the field trip
=72*37/9
=8*37
=296 students